OpenStudy (anonymous):

Awarding Fan and Medal for anybody who helps me w. 4 questions

OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-3, -1) with a slope of 4. y - 1 = 4(x - 3) y + 1 = 4(x + 3) y - 3 = 4(x + 1) y + 3 = 4(x - 1)

OpenStudy (anonymous):

If f(x) = 2x2 + 2, find f(5). 102 52 22 -1.5

OpenStudy (anonymous):

Choose the equation of the line passing through the point (2, 6) and parallel to y = -3x - 4. y = 3x - 12 y = -3x y = -3x + 12 y = 3x

OpenStudy (anonymous):

@SyedMohammed98

syed98 (syedmohammed98):

it's parallel so the gradient is same i:e -3x and use the formula